OpenStudy (anonymous):

what is the sum of the rational expressions below 2x+3/3x (plus) x/x+1

OpenStudy (anonymous):

\[\frac{2x+3}{3x}+\frac{x}{x+1}=\frac{(x+1)(2x+3)+3x\times x}{3x(x+1)}\]

OpenStudy (anonymous):

now some algebra to multiply out and combine like terms. but that is the basic idea. denominator is \[3x^2+3x\] numerator is \[2x^2+3x+2x+3+3x^2\] \[=5x^2+5x+3\] so "final answer is \[\frac{5x^2+5x+3}{3x^2+3x}\] unless i made an algebra mistake